OpenStudy (anonymous):

how to integrate cos^3(x)cos(2x)?

OpenStudy (zarkon):

\[\cos(2x)=2\cos^2(x)-1\]

OpenStudy (anonymous):

@Zarkon better way would be converting cos^3(x) to cos 3x

OpenStudy (zarkon):

but \[\cos^3(x)\ne\cos(3x)\]

OpenStudy (anonymous):

\[\cos^3x=\frac{\cos3x+3cosx}{4}\] It is lot easier to convert compound angles to integrable forms than powers of trigonometric exponents

OpenStudy (anonymous):

do you have general form?

OpenStudy (anonymous):

some one along the line forgot to invent genral formulas for integration